OpenStudy (anonymous):

Use any method to simplify the complex fraction: k+1/40k over 4k-1/32k ( the answer should look something like 4(k+4) over 3(4k-3)

OpenStudy (saifoo.khan):

\[\Large \frac{k+\frac{1}{40}k}{4k-\frac{1}{32}k}\]Like that?

OpenStudy (anonymous):

no it is k+1 over 40k and take all of that over 4k-1 over 32k

OpenStudy (anonymous):